# Weak continuous measurements require more work than strong ones

### v2

@@ VI Analysis of a single measurement: efficiency, strength and work
In the absence of coarse-graining, and for a pure ancilla initial state, the measure of efficiency $\eta$, defined in Eq. (21), is equal to 1  [^1] This remains true for any model as those conditions imply that the conditional post-measurement state is captured by a single measurement operator $\hat{\rho}_{S|r}\propto\hat{M}_{r}\hat{\rho}_{S}(0)\hat{M}_{r}^{\dagger}$. We analyze here the impact of coarse-graining by analyzing the case where the measurement result takes two values, corresponding to no excitation and any number of excitations in the ancilla, respectively (e.g. as a photodiode would do). This scheme is practically motivated for weak measurements, as a weak (short) qubit-cavity interaction makes high numbers of excitations very unlikely. This is implemented by two projectors $\hat{\Pi}_{r=e}={\left|{0}\right\rangle}{\left\langle{0}\right|}$ and $\hat{\Pi}_{r=g}=\sum_{n>0}{\left|{n}\right\rangle}{\left\langle{n}\right|}$. In this case, the efficiency measure $\eta$ depends on $\bar{\alpha}$ and $\epsilon$. As shown in Fig. 4, efficiency peaks when $\epsilon$ and $\bar{\alpha}$ are close to each other, as the conditional field displacement associated with the qubit being in the excited state becomes negligible with respect to the one associated with the ground state (and finding excitations in the cavity is strongly correlated to the qubit being in the ground state). Specifically, when $\bar{\alpha}=\epsilon$, the efficiency reaches 1. At this point, the measurement is extremely asymmetric as outcome $r=1$ is associated with the qubit excited state with certainty, while $r=0$ only provides partial information about the qubit state. As a consequence, $p(r=e)=p_{e}+p_{g}e^{-4|\bar{\alpha}|^{2}}\neq p_{e}(0)$, meaning that the measurement is biased unless $|\bar{\alpha}|\gtrsim 1$ [26].
@@
where the excitation number distribution associated with the excited state is a Kronecker delta and has zero entropy. For large $\bar{\alpha}$ such that the excitation number distribution associated with the ground state is peaked around $n\leavevmode\nobreak\ |\bar{\alpha}|^{2}\gg 1$, the Shannon entropy of the total excitation number distribution scales like the sum of the Shannon entropies of the two conditional distributions $p(n|e)$ and $p(n|g)$, and therefore drops when $\bar{\alpha}\simeq\epsilon$ (see also Appendix D for a more general discussion of this behavior.). This effect is not visible for $\bar{\alpha}\ll 1$ as $p(0|g)$ becomes non-negligible. In this protocol, the work bound clearly depends on the measurement strength. We note, however, that when using an ancilla with a bounded Hilbert space, such as a qubit, the work cost for a single measurement can be independent of the strength for specific initial system states [37].
@@
As shown in Fig. 9, this operation indeed increases the classical mutual information between the ancilla excitation number $n$ and the value of $\sigma_{z}$. Maximal values $\eta_{\sigma_{z}:r}=0.977$ are obtained for $\theta=\frac{\pi}{4}+m\frac{\pi}{2}$, where $m$ is a natural number. Further improvement might be obtained owing to a global unitary on the ancilla.
@@ VII Concatenated weak measurement versus strong measurement
In this section, we compare the work cost associated with a strong measurement obtained by concatenating a long sequence of weak measurements, versus a single iteration of the measurement procedure of section IV with large strength. When the duration of each weak measurement involved in the sequence goes to zero, a so-called weak continuous measurement is obtained [28]. We assume that, at each iteration, the whole process in Fig. 1 is performed, including the reset of the ancilla and classical memory. The system state is incremented according to Eq. (18), with $M_{n}$ being infinitesimally close to $\mathds{1}$. This cycle continues until all relevant information is extracted from the system.
From the analysis of a single measurement in previous section, the minimum work cost occurs at $\bar{\alpha}=\epsilon$ (see Fig. 6). We therefore assume this equality to hold hereafter, and analyze results in terms of $\bar{\alpha}$. We plot the work bound as a function of $\bar{\alpha}=\epsilon$ in Fig. 10, showing that it increases sub-linearly with $\bar{\alpha}$. This behavior provides a first indication that $N\gg 1$ weak measurements of strength $\xi_{\text{weak}}\ll 1$ should have a larger work cost than a single measurement of maximal strength $\xi_{\text{strong}}\sim N\xi_{\text{weak}}\sim\log 2$. We formalize this result in the remainder of this section, showing that a sequence of weak measurement costs more work than a direct implementation of the resulting strong measurement, and that this statement apply also to the other measurement figure of merits.
@@
To this goal, we analyze the behavior of a sequence weak measurements as a single stronger measurement. We thus need a rule to concatenate the $N$ measurement outcomes $\vec{r}=(r_{1},....,r_{N})$ obtained in the elementary measurements of the sequence into a single result $R=e,g$, and derive the compute qubit state update. Due to the asymmetry (and resulting bias) of the measurement associated with our example (see section VI), collapsing towards state ${\left|{e}\right\rangle}$ and ${\left|{g}\right\rangle}$ typically requires a different number of weak measurements, with relative weights depending on the parameters $\bar{\alpha}$ and $\epsilon$. As a consequence, combining the weak measurement results of a sequence $\vec{r}$ into a single result $R$ is not trivial. We therefore adopt a Bayesian systematic approach. Namely, we assign to $R$ the eigenstate (ground or excited) with highest probability, given the sequence $(r_{1},....,r_{N})$, as computed from Eq. (18) via
@@
We now analyze the figure of merits for the measurement emerging from such concatenation procedure. They are plotted in Fig. 11 as a function of the number $N$ of iterations in the sequence.
Figure 11: Strength ($\xi$), detection efficiency ($\eta$), normalized classical mutual information ($\eta_{X:r}$), and product $\xi\eta\eta_{X:r}$ as a function of the number of iterations ($N$) for different values of average displacement ($\bar{\alpha}$). The sudden jump in detection efficiency and normalized classical mutual information arises due to the concatenation rule. This occurs when an additional outcome combination sequence satisfies $p_{e|\vec{r}}>p_{g|\vec{r}}$. The qubit was initialized in the reference state ${\left|{\text{ref}}\right\rangle}=\frac{1}{\sqrt{2}}\left({\left|{g}\right\rangle}+{\left|{e}\right\rangle}\right)$.
@@
The detection efficiency measure $\eta$ is plotted in the second panel. While a single measurement has maximal detection efficiency ($\eta=1$) in the absence of coarse-graining, repeated protocol applications reduce the purity of the conditional ancilla states, making efficiency dependent on $\bar{\alpha}$ and $\epsilon$. This can be intuitively understood noting that the bayesian procedure to combine the sequence results into a two-valued outcome is a kind of coarse-graining. For sufficient large values of $\bar{\alpha}\gtrsim\epsilon$, the efficiency decreases to a minimum values over a few hundreds of repetitions, before converging to $1$ at large number of repetitions (it remains all along at exactly $1$ for $\bar{\alpha}=\epsilon$). For $\bar{\alpha}\ll\epsilon$, the efficiency drops and convergence is not ensured.
@@
We end this section by comparing the work bound $W_{1}$ associated with a single strong measurement, characterized by $\epsilon_{\text{strong}}=\bar{\alpha}_{\text{strong}}=1.5$ (achieving all three figures or merit to be larger or equal to $0.999$), versus the total work $W_{N\epsilon}$ associated to a sequence of $N$ weak measurements characterized by $\epsilon=\bar{\alpha}\ll 1$. The value of $N$ in the comparison is chosen such that resulting measurement achieves the same performance, that is all three figures of merits larger or equal to $0.999$. As shown in Fig. 12 the work cost of the measurement sequence (blue line) is typically hundred times higher than the single strong measurement (orange line), and this remains true for a large range of values of $\epsilon=\bar{\alpha}$.
@@
As we argue below, this trend is actually very general. As detailed in Appendix D, for any protocol where dephasing plays the role of converting quantum states into classical ones – a mechanism that aligns with the framework of quantum Darwinism [10] –, and largely regardless of the specific model, initial ancilla state, and many other details, we show that the work cost for $N$ concatenated weak measurements is significantly higher than that of a single strong, efficient measurement achieving similar performance. Our proof is based on phenomenological assumptions about the functional form of the ancilla’s state distribution $p_{n}$ that must be satisfied for a strong and weak measurement, respectively.
@@
[…]
@@ VIII Conclusions
This article examines a minimal setup for modeling non-ideal quantum measurements, consisting of an interacting system of interest and an ancilla. After the interaction, the ancila undergoes dephasing, transforming quantum data into classical outputs, simulating the objectification process. The ancilla is read out, and the data is stored in a classical memory. Finally, both the ancilla and the classical memory are reset to restart the measurement protocol.
@@ Appendix A Dissipation energetic calculations
In this appendix, we calculate the work bound for the scenario where dephasing in our model is replaced by dissipation. Unlike the dephasing model, driving and action of the bath can happen simultaneously [25]. We will analyze the work cost of the protocol step by step.
@@ Appendix B Dynamical evolution
[…]
@@
$\frac{\partial\hat{\rho}_{I}}{\partial t^{\prime}}=-i\left[\hat{V}_{I}\left(t^{\prime}\right),\hat{\rho}_{I}\left(t{\color[rgb]{0.41,0.05,0.68}{{}^{\prime}}}\right)\right],$ (42)
@@
[…]
@@
[…]
@@
[…]
@@
$\hat{U}_{I}\left(\Delta t\right)$ $=$ $\lim_{n\to\infty}\text{exp}\left[\sum_{l=1}^{n-1}i\sin{\frac{\omega_{a}\Delta tl}{n}}\sum_{b=1}^{n-l+1}\prod_{c=0}^{b}\mu\left(\frac{c\Delta t}{n}\right)\frac{\Delta t}{n}\right]$
@@
[…]
@@ Appendix C Initial thermal ancilla state
In this appendix, we examine the performance of the qubit measurement when the ancilla is initialized in a thermal state of finite temperature. The protocol, as described in section IV, is modified by replacing the initial ancilla vacuum state with a thermal state characterized by $\beta\omega_{B}=3$, where $\hbar=1$. This value was chosen for being of the correct order of magnitude to describe, e.g. superconducting qubit readout via their coupling to microwave cavities [38]. After applying the protocol, we evaluate the resulting performance metrics and analyze the associated work bound.
@@
In Fig. 16, we plot the work bound of the protocol, as given by Eq. 11. For the same values of $\bar{\alpha}$ and $\epsilon$, the work bound is higher when the initial ancilla state is thermal compared with the vacuum. This result is in agreement with observations in Ref.[25] that the cost increases with the amount of extracted information and the measurement efficiency, which both decrease with the temperature. Indeed, when starting with a thermal state, we have:
@@ D.1 Scaling of the work cost
In the most simple case, the interaction only affects the average of the pointer distributions, that is $\sigma_{k}(\epsilon)=\sigma_{0}$. As the Shannon entropy of a single-peak distribution typically scales as the log of its standard deviation, the work cost interpolates between two plateaus $H_{0}$ and $H_{0}+H(p_{k})$ when $\epsilon$ is varied from 00 to $\infty$. Importantly, in this case, the work cost saturates at very large $\epsilon$.
@@
In the specific example treated in this article, we can identify the functions $G_{u_{k}(\epsilon)}^{\sigma_{k}(\epsilon)}(n)\equiv e^{-|\alpha_{k}|^{2}}\frac{|\alpha_{k}|^{2n}}{n!}$ with $\alpha_{e}=\bar{\alpha}-\epsilon$ and $\alpha_{g}=\bar{\alpha}+\epsilon$. Thus, $u_{e/g}(\epsilon)=|\bar{\alpha}\mp\epsilon|^{2}$ and $\sigma_{e/g}(\epsilon)=|\bar{\alpha}\mp\epsilon|$. For a fixed $\bar{\alpha}$, when increasing $\epsilon$ from 00 to $\infty$, we can identify three regimes. For $\epsilon\ll 1$, $H(p_{n})$ is of the order of the entropy of the Shannon entropy of the Poisson distribution $e^{-|\bar{\alpha}|^{2}}\frac{|\bar{\alpha}|^{2n}}{n!}$. When $\epsilon\gg\bar{\alpha}$, $H(p_{n})$ is of the order of $H(p_{k})+H\left(e^{-|\epsilon|^{2}}\frac{|\epsilon|^{2n}}{n!}\right)$, which increases with $\epsilon$ due to the second term. In between, the variance $\sigma_{e}$ turns out to have a non monotonic behavior. In particular, when $\epsilon\simeq\bar{\alpha}$, the variance of the distribution associated to $k=e$ vanishes, and consequently so does $H(G_{u_{e}(\epsilon)}^{\sigma_{e}(\epsilon)}(n))$. This property may lead to a non-monotonic behavior for the work cost for $\epsilon\in[0,\bar{\alpha}]$. For instance, if $\bar{\alpha}\gtrsim 1$, the two distributions $G_{u_{k}(\epsilon)}^{\sigma_{k}(\epsilon)}(n)$ are already orthogonal when $\epsilon=\bar{\alpha}$, and $H(p_{n})\sim H(p_{k})+p_{g}H\left(e^{-|2\bar{\alpha}|^{2}}\frac{|2\bar{\alpha}|^{2n}}{n!}\right)$. Moreover, $H\left(e^{-|x|^{2}}\frac{|x|^{2n}}{n!}\right)$ is sublinear in $x$ as long as $x>1$. Those facts can lead to situations where the minimum work cost can decrease with $\epsilon$ in the interval $[0,\bar{\alpha}]$, or even reach its overall minimum at $\epsilon=\bar{\alpha}$ (see Fig. 6).

### v3

@@ VI Analysis of a single measurement: efficiency, strength and work
In the absence of coarse-graining, and for a pure ancilla initial state, the measure of efficiency $\eta$, defined in Eq. (22), is equal to 1  [^1] This remains true for any model as those conditions imply that the conditional post-measurement state is captured by a single measurement operator $\hat{\rho}_{S|r}\propto\hat{M}_{r}\hat{\rho}_{S}(0)\hat{M}_{r}^{\dagger}$. We analyze here the impact of coarse-graining by analyzing the case where the measurement result takes two values, corresponding to no excitation and any number of excitations in the ancilla, respectively (e.g. as a photodiode would do). This scheme is practically motivated for weak measurements, as a weak (short) qubit-cavity interaction makes high numbers of excitations very unlikely. This is implemented by two projectors $\hat{\Pi}_{r=e}={\left|{0}\right\rangle}{\left\langle{0}\right|}$ and $\hat{\Pi}_{r=g}=\sum_{n>0}{\left|{n}\right\rangle}{\left\langle{n}\right|}$. In this case, the efficiency measure $\eta$ depends on $\bar{\alpha}$ and $\epsilon$. As shown in Fig. 5, efficiency peaks when $\epsilon$ and $\bar{\alpha}$ are close to each other, as the conditional field displacement associated with the qubit being in the excited state becomes negligible with respect to the one associated with the ground state (and finding excitations in the cavity is strongly correlated to the qubit being in the ground state). Specifically, when $\bar{\alpha}=\epsilon$, the efficiency reaches 1. At this point, the measurement is extremely asymmetric as outcome $r=1$ is associated with the qubit excited state with certainty, while $r=0$ only provides partial information about the qubit state. As a consequence, $p(r=e)=p_{e}+p_{g}e^{-4|\bar{\alpha}|^{2}}\neq p_{e}(0)$, meaning that the measurement is biased unless $|\bar{\alpha}|\gtrsim 1$ [guryanova2020ideal].
@@
where the excitation number distribution associated with the excited state is a Kronecker delta and has zero entropy. For large $\bar{\alpha}$ such that the excitation number distribution associated with the ground state is peaked around $n\penalty 10000\ |\bar{\alpha}|^{2}\gg 1$, the Shannon entropy of the total excitation number distribution scales like the sum of the Shannon entropies of the two conditional distributions $p(n|e)$ and $p(n|g)$, and therefore drops when $\bar{\alpha}\simeq\epsilon$ (see also Appendix D for a more general discussion of this behavior). This effect is not visible for $\bar{\alpha}\ll 1$ as $p(0|g)$ becomes non-negligible. In this protocol, the work bound clearly depends on the measurement strength. We note, however, that when using an ancilla with a bounded Hilbert space, such as a qubit, the work cost for a single measurement can be independent of the strength for specific initial system states [mancino2018entropic].
@@
As shown in Fig. 10, this operation indeed increases the classical mutual information between the ancilla excitation number $n$ and the value of $\hat{\sigma}_{z}$. Maximal values $\eta_{\sigma_{z}:r}=0.977$ are obtained for $\theta=\frac{\pi}{4}+m\frac{\pi}{2}$, where $m$ is a natural number. Further improvement might be obtained owing to a global unitary on the ancilla.
@@ VII Concatenated weak measurement versus strong measurement
In this section, we compare the work cost associated with a strong measurement obtained by concatenating a long sequence of weak measurements, versus a single iteration of the measurement procedure of section IV with large strength. When the duration of each weak measurement involved in the sequence goes to zero, a so-called weak continuous measurement is obtained [jacobs2006straightforward]. We assume that, at each iteration, the whole process in Fig. 1 is performed, including the reset of the ancilla and classical memory. The system state is incremented according to Eq. (19), with $\hat{M}_{n}$ being infinitesimally close to $\mathds{1}$. This cycle continues until all relevant information is extracted from the system.
From the analysis of a single measurement in the previous section, the minimum work cost occurs at $\bar{\alpha}=\epsilon$ (see Fig. 7). We therefore assume this equality to hold hereafter, and analyze results in terms of $\bar{\alpha}$. We plot the work bound as a function of $\bar{\alpha}=\epsilon$ in Fig. 11, showing that it increases sub-linearly with $\bar{\alpha}$. This behavior provides a first indication that $N\gg 1$ weak measurements of strength $\xi_{\text{weak}}\ll 1$ should have a larger work cost than a single measurement of maximal strength $\xi_{\text{strong}}\sim N\xi_{\text{weak}}\sim\log 2$. We formalize this result in the remainder of this section, showing that a sequence of weak measurement costs more work than a direct implementation of the resulting strong measurement, and that this statement also applies to the other measurement figures of merit.
@@
To this goal, we analyze the behavior of a sequence of weak measurements as a single stronger measurement. We thus need a rule to concatenate the $N$ measurement outcomes $\vec{r}=(r_{1},....,r_{N})$ obtained in the elementary measurements of the sequence into a single result $R=e,g$, and derive the compute qubit state update. Due to the asymmetry (and resulting bias) of the measurement associated with our example (see section VI), collapsing towards state $\left|{e}\right\rangle$} and $\left|{g}\right\rangle$} typically requires a different number of weak measurements, with relative weights depending on the parameters $\bar{\alpha}$ and $\epsilon$. As a consequence, combining the weak measurement results of a sequence $\vec{r}$ into a single result $R$ is not trivial. We therefore adopt a Bayesian systematic approach. Namely, we assign to $R$ the eigenstate (ground or excited) with highest probability, given the sequence $(r_{1},....,r_{N})$, as computed from Eq. (19) via
@@
We now analyze the figures of merit for the measurement emerging from such concatenation procedure. They are plotted in Fig. 12 as a function of the number $N$ of iterations in the sequence.
Figure 12: Strength ($\xi$), detection efficiency ($\eta$), normalized classical mutual information ($\eta_{X:r}$), and product $\xi\eta\eta_{X:r}$ as a function of the number of iterations (NN) for different values of average displacement ($\bar{\alpha}$). The sudden jump in detection efficiency and normalized classical mutual information arises due to the concatenation rule. This occurs when an additional outcome combination sequence satisfies $p_{e|\vec{r}}>p_{g|\vec{r}}$. The qubit was initialized in the reference state ${\left|{\text{ref}}\right\rangle}=\frac{1}{\sqrt{2}}\left({\left|{g}\right\rangle}+{\left|{e}\right\rangle}\right)$.
@@
The detection efficiency measure $\eta$ is plotted in the second panel. While a single measurement has maximal detection efficiency ($\eta=1$) in the absence of coarse-graining, repeated protocol applications reduce the purity of the conditional ancilla states, making efficiency dependent on $\bar{\alpha}$ and $\epsilon$. This can be intuitively understood noting that the Bayesian procedure to combine the sequence results into a two-valued outcome is a kind of coarse-graining. For sufficient large values of $\bar{\alpha}\gtrsim\epsilon$, the efficiency decreases to minimum values over a few hundreds of repetitions, before converging to $1$ at large number of repetitions (it remains all along at exactly $1$ for $\bar{\alpha}=\epsilon$). For $\bar{\alpha}\ll\epsilon$, the efficiency drops and convergence is not ensured.
@@
We end this section by comparing the work bound $W_{1}$ associated with a single strong measurement, characterized by $\epsilon_{\text{strong}}=\bar{\alpha}_{\text{strong}}=1.5$ (achieving all three figures or merit to be larger or equal to $0.999$), versus the total work $W_{N\epsilon}$ associated to a sequence of $N$ weak measurements characterized by $\epsilon=\bar{\alpha}\ll 1$. The value of $N$ in the comparison is chosen such that the resulting measurement achieves the same performance, that is all three figures of merits larger or equal to $0.999$. As shown in Fig. 13 the work cost of the measurement sequence (blue line) is typically hundred times higher than the single strong measurement (orange line), and this remains true for a large range of values of $\epsilon=\bar{\alpha}$.
@@
As we argue below, this trend is actually very general. As detailed in Appendix D, for any protocol where dephasing plays the role of converting quantum states into classical ones – a mechanism that aligns with the framework of quantum Darwinism [zurek2009quantum] –, and largely regardless of the specific model, initial ancilla state, and many other details, we show that the work cost for $N$ concatenated weak measurements is significantly higher than that of a single strong, efficient measurement achieving similar performance. Our proof is based on phenomenological assumptions about the functional form of the ancilla’s state distribution $p_{n}$ that must be satisfied for a strong and weak measurement, respectively.
@@
Finally, we note that those results are in agreement with the experimental observations of [mancino2018entropic], where the cost of a single qubit measurement (more precisely, its entropy production) is shown to be independent of the measurement strength for their protocol, when starting in the same reference state $\left|{\text{ref}}\right\rangle$}. As a consequence, the work cost for a sequence of $N\penalty 10000\ 1/\epsilon_{\text{weak}}^{2}\gg 1$ weak measurements is expected to be $N$ times larger.
@@ VIII Conclusions
This article examines a minimal setup for modeling non-ideal quantum measurements, consisting of an interacting system of interest and an ancilla. After the interaction, the ancilla undergoes dephasing through different channels, transforming quantum data into classical outputs in the form of a SBS state, simulating the objectification process. The ancilla is read out, and the data is stored in a classical memory. Finally, both the ancilla and the classical memory are reset to restart the measurement protocol.
@@ Appendix A Dissipation energetic calculations
In this appendix, we calculate the work bound for the scenario where dephasing in our model is replaced by dissipation. Unlike the dephasing model, driving and action of the bath can happen simultaneously [latune2024thermodynamically]. We will analyze the work cost of the protocol step by step.
@@ Appendix B Dynamical evolution
$\hat{\rho}_{I}\left(t_{1}\right)=\hat{U}_{0}^{\dagger}\left(\Delta t\right)\hat{\rho}_{S}\left(t_{0}\right)\hat{U}_{0}\left(\Delta t\right)\hskip 28.45274pt\hat{V}_{I}\left(\Delta t\right)=\hat{U}_{0}\left(\Delta t\right)^{\dagger}\hat{V}\left(\Delta t\right)\hat{U}_{0}\left(\Delta t\right),$ (42)
@@
$\frac{\partial\hat{\rho}_{I}}{\partial t^{\prime}}=-i\left[\hat{V}_{I}\left(t^{\prime}\right),\hat{\rho}_{I}\left(t{\color[rgb]{0,0,0}{{}^{\prime}}}\right)\right],$ (43)
@@
$\hat{\rho}_{I}\left(t_{1}\right)=\hat{\rho}_{I}\left(t_{0}\right)-i\int_{t_{0}}^{t_{1}}\left[\hat{V}_{I}\left(t^{\prime}\right),\hat{\rho}_{I}\left(t^{\prime}\right)\right]dt^{\prime}.$ (44)
@@
$\hat{U}_{I}\left(\Delta t\right)=\mathcal{T}\left(\text{exp}\left(-i\int_{t_{0}}^{t_{1}}\hat{V}_{I}\left(t^{\prime}\right)dt^{\prime}\right)\right).$ (48)
@@
$U_{I,g}\left(\Delta t\right)=\mathcal{T}\left(\text{exp}\left(-i\int_{t_{0}}^{t_{1}}\mu\left(t^{\prime}\right)\left(\hat{a}^{\dagger}e^{i\omega_{a}t^{\prime}}+\hat{a}e^{-i\omega_{a}t^{\prime}}\right)dt^{\prime}\right)\right),$ (49)
@@
[…]
@@
$\alpha=-i\int_{t_{0}}^{t_{1}}\mu\left(t^{\prime}\right)e^{i\omega_{a}t^{\prime}}dt^{\prime}.$ (54)
@@ Appendix C Initial thermal ancilla state
In this appendix, we examine the performance of the qubit measurement when the ancilla is initialized in a thermal state of finite temperature. The protocol, as described in section IV, is modified by replacing the initial ancilla vacuum state with a thermal state characterized by $\beta\omega_{B}=3$, where $\hbar=1$. This value was chosen for being of the correct order of magnitude to describe, e.g. superconducting qubit readout via their coupling to microwave cavities [ficheux2018dynamics]. After applying the protocol, we evaluate the resulting performance metrics and analyze the associated work bound.
@@
In Fig. 17, we plot the work bound of the protocol, as given by Eq. 12. For the same values of $\bar{\alpha}$ and $\epsilon$, the work bound is higher when the initial ancilla state is thermal compared with the vacuum. This result is in agreement with observations in Ref.[latune2024thermodynamically] that the cost increases with the amount of extracted information and the measurement efficiency, which both decrease with the temperature. Indeed, when starting with a thermal state, we have:
@@ D.1 Scaling of the work cost
In the most simple case, the interaction only affects the average of the pointer distributions, that is $\sigma_{k}(\epsilon)=\sigma_{0}$. As the Shannon entropy of a single-peak distribution typically scales as the log of its standard deviation, the work cost interpolates between two plateaus $H_{0}$ and $H_{0}+H(p_{k})$ when $\epsilon$ is varied from 0 to $\infty$. Importantly, in this case, the work cost saturates at very large $\epsilon$.
@@
In the specific example treated in this article, we can identify the functions $G_{u_{k}(\epsilon)}^{\sigma_{k}(\epsilon)}(n)\equiv e^{-|\alpha_{k}|^{2}}\frac{|\alpha_{k}|^{2n}}{n!}$ with $\alpha_{e}=\bar{\alpha}-\epsilon$ and $\alpha_{g}=\bar{\alpha}+\epsilon$. Thus, $u_{e/g}(\epsilon)=|\bar{\alpha}\mp\epsilon|^{2}$ and $\sigma_{e/g}(\epsilon)=|\bar{\alpha}\mp\epsilon|$. For a fixed $\bar{\alpha}$, when increasing $\epsilon$ from 0 to $\infty$, we can identify three regimes. For $\epsilon\ll 1$, $H(p_{n})$ is of the order of the entropy of the Shannon entropy of the Poisson distribution $e^{-|\bar{\alpha}|^{2}}\frac{|\bar{\alpha}|^{2n}}{n!}$. When $\epsilon\gg\bar{\alpha}$, $H(p_{n})$ is of the order of $H(p_{k})+H\left(e^{-|\epsilon|^{2}}\frac{|\epsilon|^{2n}}{n!}\right)$, which increases with $\epsilon$ due to the second term. In between, the variance $\sigma_{e}$ turns out to have a non monotonic behavior. In particular, when $\epsilon\simeq\bar{\alpha}$, the variance of the distribution associated to $k=e$ vanishes, and consequently so does $H(G_{u_{e}(\epsilon)}^{\sigma_{e}(\epsilon)}(n))$. This property may lead to a non-monotonic behavior for the work cost for $\epsilon\in[0,\bar{\alpha}]$. For instance, if $\bar{\alpha}\gtrsim 1$, the two distributions $G_{u_{k}(\epsilon)}^{\sigma_{k}(\epsilon)}(n)$ are already orthogonal when $\epsilon=\bar{\alpha}$, and $H(p_{n})\sim H(p_{k})+p_{g}H\left(e^{-|2\bar{\alpha}|^{2}}\frac{|2\bar{\alpha}|^{2n}}{n!}\right)$. Moreover, $H\left(e^{-|x|^{2}}\frac{|x|^{2n}}{n!}\right)$ is sublinear in $x$ as long as $x>1$. Those facts can lead to situations where the minimum work cost can decrease with $\epsilon$ in the interval $0,\bar{\alpha}]$, or even reach its overall minimum at $\epsilon=\bar{\alpha}$ (see Fig. 7).